OpenStudy (simplymarie_x):

Explain the difference between using the sine ratio to solve for a missing angle in a right triangle versus using the cosecant ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (johnweldon1993):

Hint* \[\large sin(\theta) = \frac{opposite}{hypotenuse}\] Whereas \[\large csc(\theta) = \frac{1}{sin(\theta)} = \frac{hypotenuse}{opposite}\]